### v4

# Unbiased Region-Language Alignment for Open-Vocabulary Dense Prediction

Abstract
Pre-trained vision-language models (VLMs), such as CLIP, have demonstrated impressive zero-shot recognition capability, but still underperform in dense prediction tasks. Self-distillation recently is emerging as a promising approach for fine-tuning VLMs to better adapt to local regions without requiring extensive annotations. However, previous state-of-the-art approaches often suffer from significant ‘foreground bias’, where models tend to wrongly identify background regions as foreground objects. To alleviate this issue, we propose DenseVLM, a framework designed to learn unbiased region-language alignment from powerful pre-trained VLM representations. DenseVLM leverages the pre-trained VLM to retrieve categories for unlabeled regions and then decouples the interference between foreground and background features. This separation ensures accurate region-category alignment while maintaining semantic distinctions during training. We show that DenseVLM can directly replace the original VLM in open-vocabulary object detection and image segmentation methods, leading to notable performance improvements. Furthermore, it exhibits promising zero-shot scalability when training on more extensive and diverse datasets. Our code is publicly available https://github.com/HVision-NKU/DenseVLM.
22footnotetext: Corresponding Author. Email: houqb@nankai.edu.cn
1 Introduction
Open-vocabulary dense prediction, primarily including object detection [Zareian_2021_CVPR, gu2021open, minderer2022simple, kuo2022f, jia2024mssd] and image segmentation [openseg, odise, ovseg, san, catseg, li2024cascade], aims to identify regions or dense visual concepts of arbitrary categories as described by the text. Benefiting from the powerful pre-trained Vision-Language Models (VLMs), recent open-vocabulary approaches [maskclip1, Kim_2023_CVPR, san] for dense prediction have achieved great progress.
[…]
Figure 1: Illustration of foreground bias. Previous methods [sun2023eva, zhong2022regionclip, wu2023clipself] often produce similar foreground predictions for background regions, our approach effectively alleviates this issue.
Popular VLMs, such as CLIP [CLIP] and EVA-CLIP [sun2023eva], have exhibited remarkable zero-shot recognition abilities for global image understanding [huang2024class, huang2025mindgappreservingcompensating, li2024advancing, li2024promptkd, fang2025aligning]. However, these models expose notable limitations in the understanding of local visual semantics, particularly in localizing and identifying small objects and background stuff [zhong2022regionclip, maskclip1]. This limitation arises from the training manner of VLMs that align images with global text while neglecting the correspondences between local image regions and their specific text descriptions. To alleviate this issue, some studies use region-text or pseudo region-text pairs [zhong2022regionclip, li2022grounded, zhang2022glipv2, liu2023grounding] but these methods are limited by the high annotation cost and lack scalability for open-world scenes. In contrast, self-supervised approaches, such as CLIPSelf [wu2023clipself] and MaskEmbed [covert2024locality], align region semantics using cropped image representations or reconstruct masked patch embeddings, respectively. These self-distillation approaches, which do not rely on annotated data, offer flexibility and scalability across a variety of datasets.
[…]
Figure 2: Comparison of different VLMs. Unlike existing methods using (a) image-text contrastive learning [CLIP], (b) region-text contrastive learning [zhong2022regionclip] or (c) self-distillation [wu2023clipself], our method leverages powerful model representations for region-language alignment.
Despite the advances, previous VLMs [CLIP, sun2023eva, zhong2022regionclip] pre-trained on image-text and region-text pairs tend to overemphasize the foreground objects at the expense of the background context. This disproportionate focus results in a pronounced predilection for foreground object recognition and causes models to incorrectly associate background regions with foreground labels. Consequently, in dense prediction tasks, these VLMs often misclassify background regions to co-occurring foreground classes—a phenomenon we term as ‘foreground bias.’ To illustrate this issue, we present an example comparing the classification results of popular VLMs [sun2023eva, zhong2022regionclip, wu2023clipself], where features are extracted from regions using ground-truth masks. As illustrated in Fig. 1, these models tend to confuse ‘sky’ (a background class) with ‘building’ (a foreground object).
To address this bias, we propose aligning foreground and background regions separately, ensuring explicit semantic separation through distinct category sets. To achieve this, we introduce DenseVLM, an end-to-end framework designed for unbiased region-language alignment, as shown in Fig. 2(d). Specifically, for unlabeled regions, we leverage a powerful pre-trained VLM that has learned robust semantic features from diverse data to retrieve relevant categories without relying on paired data or self-distillation [sun2023eva, zhong2022regionclip, wu2023clipself]. To ensure semantic diversity, DenseVLM incorporates a broad and comprehensive set of categories derived from large-scale datasets or predicted by generative models [Yue_2024_CVPR]. A key feature of DenseVLM is its ability to classify regions as either foreground or background based on predefined sets for each category. This classification way enables the decoupling of region features, reducing interactions between foreground and background. By decoupling these features, DenseVLM achieves independent region alignment of their respective categories while maintaining a semantic separation during training. Furthermore, DenseVLM improves both efficiency and performance by directly extracting region features from the dense features of VLMs, avoiding traditional image cropping [zhu2024mrovseg, wu2023clipself].
We evaluate the effectiveness of DenseVLM on several open-vocabulary benchmarks [wu2023clipself], including object detection (box classification) and image segmentation (thing and stuff mask recognition). DenseVLM is adaptable to various network architectures, such as ViTs [dosovitskiy2020image] and CNNs [lecun1989backpropagation], and consistently outperforms other competing methods [CLIP, sun2023eva, wu2023clipself]. Furthermore, DenseVLM has great scaling ability, showing promising performance improvement when scaling up the training set based on the SA-1B [kirillov2023segment] dataset. For downstream tasks, DenseVLM improves the two-stage models [kuo2022f] of OV-COCO [chen2015microsoft] and OV-LVIS [Gupta_2019_CVPR] in open-vocabulary object detection and achieves notable gains in open-vocabulary semantic segmentation over state-of-the-art methods such as SAN [san] and Cat-Seg [catseg]. We summarize our contributions as follows.
  • We identify the foreground bias issue in existing VLMs and propose region-text alignment by incorporating explicit semantic structuring through category guidance.
  • We propose DenseVLM, a region-language alignment framework that leverages a powerful VLM to retrieve categories for unlabeled regions and decouples foreground and background features to reduce foreground bias.
  • Extensive experiments on dense prediction benchmarks show that our DenseVLM outperforms previous methods and exhibits promising scalability.
2 Related work
Open-vocabulary dense prediction. Open-vocabulary dense prediction approaches aim to overcome the constraints of predefined categories, thereby enhancing their application in object detection [Zareian_2021_CVPR, gu2021open, minderer2022simple, kuo2022f] and image segmentation [openseg, odise, ovseg, san, li2024cascade, li2024clip]. The success of pre-trained vision-language models like CLIP [CLIP], has further inspired advancements in this area. In open-vocabulary detection, recent studies [gu2021open, Wu_2023_CVPR] exploit the CLIP models to effectively identify novel objects. Furthermore, several works [kuo2022f, CORA] construct object detectors that utilize frozen CLIP encoders, thereby reducing computational overhead while maintaining performance. For open-vocabulary segmentation, a common two-stage pipeline [maskclip, zegformer, odise, liu2023open] integrates a class-agnostic mask generator with a fixed CLIP encoder to achieve cross-modal alignment and mask classification. Recent methods also investigate the use of shared frozen CLIP with side adapter networks [xu2023san] or the adoption of single-stage frameworks [fcclip]. However, due to its training on image-text pairs, CLIP lacks precise local vision-language alignment, which is essential for dense prediction tasks. Although recent studies fine-tune CLIP for pixel-level [catseg, xie2023sed] or mask-level [jiao2024collaborative] alignment, they are constrained by the scarcity of densely labeled data.
[…]
Figure 3: Overview of DenseVLM. DenseVLM leverages the powerful VLM to retrieve categories for unlabeled regions and distinguish between foreground and background. During VLM training, DenseVLM decouples interference between features of foreground and background regions, aligning each region with its corresponding text embeddings.
Vision-language alignment at image and region levels. Pre-training visual-language models enable alignment between image and text representations [align, Vilbert, CLIP, jose2025dinov2, ma2024clip, chen2023vlp, wang2023large]. By using image-text pairs [schuhmann2022laion] as training data, methods such as CLIP [CLIP] and ALIGN [align] exhibit impressive zero-shot classification capabilities. To improve the alignment of vision-language of dense features, some training-free studies [wang2025sclip, lan2024clearclip, lan2024proxyclip, wysoczanska2024clip] attempt to modify the output layers of CLIP. Recent methods [mukhoti2023open, ranasinghe2023perceptual, naeem2024silc] align visual patches with image-level text or learn local-global correspondence via self-distillation, both requiring extensive text-image pairs.
For precise local region alignment, researchers utilize annotations from visual grounding datasets [liu2023grounding] to train models on region-text alignment. For example, some methods such as GLIP [li2022grounded, zhang2022glipv2] and Grounding DINO [liu2023grounding] learn region-language grounding from region-text pairs or generates pseudo region-text pairs like RegionCLIP [zhong2022regionclip]. In the context of open-vocabulary detection and segmentation, some works achieve dense visual and text alignment by using mask attention [maskclip, maskclip1, Kim_2023_CVPR, san] or fine-tuning CLIP [xie2023sed, catseg, jiao2023learning]. However, these methods are constrained by the high cost of annotation, making large-scale deployment challenging. To address the issue of annotated data scarcity, recent methods like CLIPSelf [wu2023clipself] use cropped images for semantic alignment, while MaskEmbed [covert2024locality] leverages masked patch embeddings for feature reconstruction. Despite these advances, the efficacy of self-distillation methods is constrained by the suboptimal performance of the teacher model and is further compromised by foreground bias. To overcome these limitations, we leverage a powerful VLM to retrieve categories for unlabeled regions while decoupling foreground and background features through textual category guidance.
3 Method
Our aim is to develop a region-language alignment model that can effectively align local visual and semantic features, thereby enhancing the performance of VLMs in open-vocabulary dense prediction tasks. To achieve this, it is crucial to alleviate the foreground bias problem prevalent in previous VLMs [CLIP, sun2023eva] that arises from training on image-text pairs. Moreover, our approach seeks to receive better performance beyond the constraints of self-distillation [wu2023clipself].
3.1 VLM’s representation
VLMs are typically designed to learn both global visual and textual representations within a shared semantic space. In ViT-based VLMs [openclip, sun2023eva], dense visual features are extracted through residual attention blocks. Following [maskclip1, wu2023clipself], we derive the dense image feature map $\mathcal{F}$ by modifying the final residual attention block, i.e., removing the SoftMax operation and incorporating mapping layers. For a set of categories $\{c^{1},\dots,c^{D}\}$, where $D$ is the total number of categories, textual descriptions are generated using a prompt-template strategy [gu2021open], like “This is a photo of the c{c} in the scene.” These prompts are then encoded into text embeddings $\mathcal{T}=\{\mathcal{T}^{1},\ldots,\mathcal{T}^{D}\}$ by the text encoder.
3.2 DenseVLM framework
We propose DenseVLM, an end-to-end region-language alignment framework designed to mitigate foreground bias. DenseVLM achieves this by precisely aligning unlabeled regions with their corresponding categories. As shown in Fig. 3, the framework consists of two key components. First, it retrieves category semantics for region features using the P-VLMs (Powerful VLMs) with frozen weights. Second, it decouples these region-language alignments into foreground and background to train U-VLMs (Unbiased VLMs) without foreground bias. In particular, DenseVLM operates without requiring any annotations, relying on diverse category semantics from large-scale datasets [caesar2018coco, ade20k] or generating category sets from images using the NXTP [Yue_2024_CVPR].
Image patches to semantic regions. To achieve region alignment, the patch-level visual features from Sec. 3.1 need to be transformed into semantic region features. We adopt a strategy similar to [wu2023clipself], dividing the dense feature map into an $m\times n$ grid of patches. Unlike this approach, we refrain from directly cropping the input image, thereby enhancing both computational efficiency and representation effectiveness. In each training iteration, $m$ and $n$ are randomly selected from the set $\{2,\cdots,M\}$, where $M$ defaults to 6, allowing for varying patch grid sizes. The semantic region representation $\{\mathcal{F}^{1},\cdots,\mathcal{F}^{m\times n}\}$ is then extracted from dense feature map $\mathcal{F}$ via pooling (RoIAlign) [he2017mask]. This patch sampling strategy effectively captures region features of both foreground objects and background scenes. However, due to the foreground bias in VLMs [CLIP, sun2023eva, wu2023clipself], we observe that background regions are usually misclassified as foreground classes, despite containing only a small proportion of foreground-related patches.
Powerful VLM for retrieval regions. Building on the dense representation extracted from VLM described in Sec. 3.1 and its ability to map image patches to semantic regions, the powerful VLM enables the extraction of both region features $\mathcal{F}_{P}=\{\mathcal{F}^{1}_{P},\cdots,\mathcal{F}^{m\times n}_{P}\}$ and text embeddings $\mathcal{T}_{P}=\{\mathcal{T}^{1}_{P},\ldots,\mathcal{T}^{D}_{P}\}$. Next, unlabeled regions are retrieved and matched with the most relevant categories by computing the cosine similarity between the region features $\mathcal{F}_{P}$ and text embeddings $\mathcal{T}_{P}$. For a specific region $k$, the cosine similarity between its features $\mathcal{F}^{k}_{P}=\mathcal{F}_{P}[k,:]$ and the text embeddings of all categories is computed as:
$\cos(\mathcal{F}^{k}_{P},\mathcal{T}^{i}_{P})=\frac{\mathcal{F}^{k}_{P}\cdot\mathcal{T}^{i}_{P}}{\lVert\mathcal{F}^{k}_{P}\rVert\lVert\mathcal{T}^{i}_{P}\rVert},\quad\forall i=1,2,\ldots,D$ (1)
where $\cdot$ denotes the dot product and $\lVert\cdot\rVert$ represents the Euclidean norm. The probability of associating this region with the categories is determined as follows:
$p^{k}(y=c|\mathcal{F}^{k}_{P},\mathcal{T}_{P})=\frac{\exp(\cos(\mathcal{F}^{k}_{P},\mathcal{T}_{P}^{c})/\tau)}{\sum_{j=1}^{D}\exp(\cos(\mathcal{F}^{k}_{P},\mathcal{T}_{P}^{j})/\tau)},$ (2)
where $\tau=0.01$ is a temperature hyperparameter.
The use of random grids to extract region features often introduces uncertainty in fully covering a single object, especially in scenarios with multiple objects. This uncertainty significantly affects the precision of the region-category alignment. To mitigate this issue, we leverage a region denoising scheme by discarding any region whose matching probability falls below a threshold $\theta$, which is set to 0.3 by default. Therefore, this retrieval and denoising process yields a more precise and reliable alignment of regions with their corresponding categories. The optimal category for each region $k$, denoted as $c^{k}$, is determined by: $c^{k}=\operatorname{argmax}(p^{k})$, where each region-category relation is represented as $(k,c^{k})$. This alignment is fundamental for the subsequent decoupled alignment process.
COCO ADE20K
Boxes Masks-T Masks-S Boxes Masks-T Masks-S
Method Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
OpenCLIP [openclip] 49.8 74.3 51.9 72.2 29.2 54.9 28.4 54.1 29.6 53.4 37.9 66.6
DFN [fang2023data] 38.3 65.0 31.0 57.0 26.4 54.9 30.6 57.9 24.2 49.9 32.2 57.7
SigLIP [siglip] 39.9 61.4 40.4 60.1 30.3 56.4 25.9 49.2 27.3 47.6 34.5 57.3
EVA-CLIP [sun2023eva] 44.3 68.7 44.7 66.0 26.2 51.9 33.0 57.6 33.9 56.2 36.2 62.3
RegionCLIP [zhong2022regionclip] 68.5 89.5 60.7 84.3 22.0 53.5 43.2 72.2 34.0 62.6 37.7 68.6
FineCLIP [jing2025fineclip] 64.7 86.1 62.5 80.9 36.9 70.3 43.9 71.2 45.5 68.6 46.0 74.8
CLIPSelf [wu2023clipself] 69.1 88.2 66.7 83.0 41.7 75.2 48.1 77.7 47.5 74.2 53.7 82.8
DenseVLM (Ours) 72.3 89.9 70.1 84.4 44.9 76.4 51.0 81.8 49.3 76.5 57.0 84.0
Table 1: Comparisons of different models on dense representation. We report the Top1 and Top5 mean accuracy on classifying boxes and panoptic masks (thing and stuff). denotes models trained on the COCO and evaluated in a zero-shot setting on the ADE20K dataset.
[…]
Figure 4: Mask accuracy comparison across categories in COCO dataset. Our method achieves notable improvements, especially in addressing foreground bias. The foreground categories are shown in black, and the background categories are highlighted in red.
Decoupled region alignment to train VLMs without foreground bias. Upon the established region-category relations, we perform alignment between the region and text representations to train the U-VLM. A straightforward approach would be to directly align region features and text embeddings of their corresponding categories while maximizing the separation of unrelated categories. However, since the U-VLM directly inherits from the P-VLM, our experimental results in Sec. 4.3 reveal that this method exacerbates foreground bias, leading to improved foreground detection but limited gains in background recognition.
To mitigate this issue, we propose a decoupled alignment strategy that separates the alignment process for foreground and background regions. Specifically, we denote the region features of the training U-VLM as $\mathcal{F}_{U}$, using the same partitioning grid as the P-VLM. The text embeddings are represented as $\mathcal{T}_{U}$. By leveraging the region-category relations $(k,c^{k})$ retrieved by the P-VLM, we establish a one-to-one mapping for the U-VLM to associate the region features and their corresponding category embeddings. To distinguish the semantic regions of foreground and background, we decouple these region-category relations into two distinct groups, following the predefined category sets: foreground objects Thing ($\mathcal{U}$) and background regions Stuff ($\mathcal{V}$). By selectively contrasting against text embeddings related to irrelevant categories, we guide the model to focus more on the relevant background regions, reducing the impact of irrelevant foreground categories. This selective contrast helps the model capture the distinctive characteristics of background regions, leading to a more accurate separation between foreground and background. The alignment process can be effectively achieved by maximizing the cosine similarity for the region features and the text embeddings. According to Eqn. (2), when $c^{k}\in\mathcal{V}$, the probability $q^{k}$ for a specific region is calculated as:
$q^{k}=\frac{\exp(\cos(\mathcal{F}^{k}_{U},\mathcal{T}_{U}^{c})/\tau)}{\sum_{j=1}^{\mathcal{U}\cup\mathcal{V}}\exp(\cos(\mathcal{F}^{k}_{U},\mathcal{T}_{U}^{j})/\tau)}.$ (3)
Similarly, when $c^{k}\in\mathcal{U}$, the probability $\tilde{q}^{k}$ is computed as:
$\tilde{q}^{k}=\begin{cases}\frac{\exp(\cos(\mathcal{F}^{k}_{U},\mathcal{T}_{U}^{c})/\tau)}{\sum_{j=1}^{\mathcal{U}}\exp(\cos(\mathcal{F}^{k}_{U},\mathcal{T}_{U}^{j})/\tau)}&\text{if}~c\in\mathcal{U},\\ 0&\text{otherwise}.\end{cases}$ (4)
End-to-end optimization. As in Eqn. (3) and Eqn. (4), we also compute $p^{k}$ and $\tilde{p}^{k}$ for P-VLM. The proposed method DenseVLM supports end-to-end training through KL-divergence optimization as follows.
$\mathcal{L}^{k}=\begin{cases}\operatorname{KL}(p^{k}||q^{k})&\text{if}~c^{k}\in\mathcal{U},\\ \operatorname{KL}(\tilde{p}^{k}||\tilde{q}^{k})&\text{otherwise}.\end{cases}$ (5)
The overall loss for each image is computed as $\mathcal{L}=\frac{1}{m^{\prime}\times n^{\prime}}\sum^{m^{\prime}\times n^{\prime}}_{k}\mathcal{L}^{k}$, where the sum is taken over all regions, excluding those removed by the region denoising scheme.
4 Experiments
VLMs Region Alignment GPU Memory Time Overhead Boxes Masks-T Masks-S
Frozen & Training Cropping Strategy (per card) (per epoch) Top1 Top5 Top1 Top5 Top1 Top5
ViT-B/16 & ViT-B/16 Images Features KD 37G 25min 69.1 88.2 66.7 83.4 41.7 75.2
ViT-L/14 & ViT-B/16 Images Features KD 39G 37min 24.2 52.4 23.4 51.1 10.1 39.1
ViT-L/14 & ViT-B/16 Images Logics KD 39G 55min 72.2 89.8 68.8 83.8 42.6 75.2
ViT-L/14 & ViT-B/16 Features DenseVLM 39G 23min 73.4 90.5 71.0 84.8 45.6 77.8
Table 2: Results and comparisons of various frameworks, including target VLMs for aligning representations, region cropping of frozen VLMs, and optimization strategies for region-language alignment, with GPU memory efficiency and time overhead. All models are trained on four A40 GPUs, with each epoch containing 118kk images.
4.1 Benchmarks
Experiment settings. To verify the effectiveness of the proposed DenseVLM, we perform experiments on dense prediction tasks using the COCO panoptic [lin2014microsoft] val2017 split and ADE20K panoptic [zhou2017scene] val split. Following CLIPSelf [wu2023clipself], we evaluate box classification using pooled box features (Boxes) and mask classification with pooled mask features, distinguishing foreground objects (Masks-T) from background content (Masks-S). This process follows a similar procedure as illustrated in Fig. 1, where we use ground truth annotations to extract local features and evaluate the local classification. Results are reported in Top-1 and Top-5 mean accuracy across all experiments.
[…]
4.2 Comparison with other VLMs
Quantitative evaluation. We perform a comprehensive quantitative evaluation of dense representations across multiple VLMs on the COCO Panoptic [lin2014microsoft] and ADE20K Panoptic [zhou2017scene] datasets. As shown in Tab. 1, while previous methods [openclip, fang2023data, siglip, sun2023eva] achieve strong zero-shot image classification performance, their performance in region recognition is notably suboptimal. For instance, EVA-CLIP [sun2023eva] achieves only a Top-1 box classification accuracy of 44.3% on COCO and 33.0% on ADE20K. Although RegionCLIP [zhong2022regionclip], which is trained on region-text pairs, shows improved performance on COCO, but fails to generalize to datasets such as ADE20K. Moreover, FineCLIP [jing2025fineclip] and CLIPSelf [wu2023clipself], which incorporate self-distillation, achieve higher Mask-T classification accuracy but perform poorly in Mask-S classification. In contrast, our method, DenseVLM, notably outperforms these models, achieving a 4.3% improvement in Top-1 Mask-T accuracy and a 3.9% improvement in Top-1 Mask-S accuracy on COCO, highlighting its effectiveness in dense prediction tasks.
Fig. 4 shows a comparison of mask accuracy across different categories, further demonstrating the effectiveness of DenseVLM in region-text alignment. Notably, our method notably improves accuracy in background categories such as ‘sky’ and ‘wall’, which can mitigate foreground bias.
[…]
Figure 5: Comparing the alignment effect of our DenseVLM with other methods through visualizations of cosine similarity maps between visual features and text embeddings.
Qualitative results. We visualize attention maps using cosine similarity for text-described object categories. As shown in Fig. 5, DenseVLM achieves more accurate and complete object localization than EVA-CLIP and CLIPSelf. Moreover, it better preserves semantic separation, reducing interference from other objects.
4.3 Ablation study
In our ablation study, DenseVLM is trained on the unlabeled images from the COCO train2017 split using $\timesA4$0 GPUs and evaluated on the val2017 split. The experiments in Sec. 4.3 utilize the ViT-B/16 model from EVA-CLIP [sun2023eva] for its superior efficiency and capacity.
Framework design exploration. As shown in Tab. 2, we conduct ablation studies to explore key design choices in DenseVLM, including target VLMs for aligning representations, region cropping, and optimization strategies with a focus on GPU memory usage and time overhead. Using a self-distillation strategy [wu2023clipself] as the baseline, we observe a significant performance drop when replacing the target model with ViT-L/14 CLIP from CLIPSelf, due to disruption in visual-language alignment. Replacing feature distillation with logit distillation improves performance but decreases training efficiency due to repeated feature extraction while introducing foreground bias issues. In contrast, DenseVLM leverages a more efficient feature-cropping strategy and a decoupled alignment framework to achieve superior performance while substantially reducing both training time and GPU memory usage.
P-Thing P-Stuff
Thing Stuff Thing Stuff Boxes Masks-T Masks-S
74.3 70.9 42.6
74.2 70.8 42.0
74.1 70.9 41.3
73.4 71.0 45.6
Table 3: Ablation study on decoupled alignment. ✓denotes that a region is separated from the categories in this set.
Ablation study on decoupled alignment. Tab. 3 analyzes the effect of the decoupled alignment strategy on DenseVLM. The retrieved category $c^{k}$ can be categorized into Thing and Stuff, referred to as ‘P-Thing’ and ‘P-Stuff’, respectively. DenseVLM leverages $c^{k}$ for regions to selectively contrast against irrelevant Thing and Stuff categories. When regions contrast against all categories, the model can better identify the foreground regions but struggles with distinguishing the background (1st row). In the fully decoupled setting, where each region contrasts against its own set of categories, performance further degrades (2nd row). The most pronounced foreground bias occurs when P-Thing contrasts against Stuff but P-Stuff does not contrast against Thing (3rd row). In DenseVLM, we adopt a balanced strategy where P-Stuff contrasts against both Thing and Stuff, while P-Thing does not contrast against Stuff, resulting in improved overall performance by mitigating bias (4th row).
[…]
Figure 6: Zero-shot comparisons of models pre-trained on datasets with three different scales. We select three training sets from the SA-1B dataset [kirillov2023segment]: 100K, 1.1M, and 5.5M seen samples and perform the zero-shot evaluation on the COCO and ADE20K benchmarks.
Ablation study on different category sets. Tab. 4 presents the effects of different category sets used for retrieval and alignment performance. When using the category set from dataset annotations, we observe that as the number of categories increases, performance consistently improves. Adding only the background category helps the model distinguish foreground differences, as seen in rows 1 and 2, and incorporating more foreground categories leads to performance gains, as seen in rows 3 and 4. These improvements can be attributed to the enhanced representational capacity afforded by a broader category set.
To address the limitations of category set annotations, we generate category sets using the generative model NXTP [Yue_2024_CVPR] based on images and use DeepSeek-R1 [guo2025deepseek] to classify foreground and background categories. As shown in Tab. 4, our model still achieves strong performance.
Boxes Masks-T Masks-S
Categories Top1 Top5 Top1 Top5 Top1 Top5
The category set is from the dataset annotation.
133 (80) 71.1 88.5 68.7 83.0 44.7 75.2
171 (80) 72.3 89.8 69.4 85.8 44.2 76.0
273 (160) 72.3 89.9 70.1 84.4 44.9 76.4
316 (204) 73.4 90.5 71.0 84.8 45.6 77.8
The category set is generated by NXTP based on the images.
210 (133) 72.6 90.2 70.4 84.6 41.5 75.7
794 (484) 72.5 90.4 70.3 84.8 44.1 76.2
Table 4: Ablation study on different category sets. The numbers in parentheses denote the number of foreground thing classes.
Ablation study on different P-VLMs. Tab. 5 shows that using ViT-B/16 as the retrieval model, our method performs well in mask classification, outperforming the P-VLM (ViT-B/16 with CLIPSelf). This improvement effectively mitigates foreground bias by enhancing the distinction between foreground and background regions. Using ViT-L/14 further improves performance across all metrics.
4.4 Zero-shot comparisons at different data scales
[…]
Boxes Masks-T Masks-S
P-VLM Top1 Top5 Top1 Top5 Top1 Top5
ViT-B/16 70.0 88.6 68.0 83.0 43.0 75.3
ViT-L/14 73.4 90.5 71.0 84.8 45.6 77.8
Table 5: Ablation study on different P-VLMs utilized for retrieving categories for unlabeled regions.
Backbones VLMs Boxes Masks-T Masks-S
ViT-L/14 CLIPSelf 75.2 73.1 44.5
ViT-B/16 OpenCLIP 49.8 51.9 29.2
ViT-B/16* CLIPSelf 67.6 64.4 44.5
ViT-B/16* DenseVLM 71.9 70.0 47.8
ViT-B/16 EVA-CLIP 44.3 44.7 26.2
ViT-B/16 CLIPSelf 69.1 66.7 41.7
ViT-B/16 DenseVLM 73.4 71.0 45.6
R50x4 OpenCLIP 59.2 50.5 39.1
R50x4 CLIPSelf 59.1 49.9 37.2
R50x4 DenseVLM 65.6 55.2 40.9
ConvNeXt-B OpenCLIP 57.5 48.0 31.1
ConvNeXt-B CLIPSelf 62.6 57.6 41.6
ConvNeXt-B DenseVLM 67.1 63.4 43.5
Table 6: Results and comparisons of different backbones. * indicates the model initialized by OpenCLIP [openclip].
Method A-150 A-847 PC-459
PACL [mukhoti2023open] 31.4 - -
OVSeg [ovseg] 24.8 7.1 11.0
MAFT [jiao2023learning] 29.1 10.1 12.6
SED [xie2023sed] 31.6 11.4 18.6
SCAN [Liu_2024_CVPR] 30.8 10.8 13.2
CAT-Seg+CLIPSelf [wu2023clipself] 29.7 10.1 -
CAT-Seg+FineCLIP [jing2025fineclip] 32.4 12.2 -
SAN [san] 27.4 10.0 13.0
SAN+DenseVLM 29.5+2.1 10.4+0.4 15.6+2.6
CAT-Seg [catseg] 31.4 11.7 18.4
CAT-Seg+DenseVLM 34.1+2.7 12.2+0.5 18.7+0.3
Table 7: Results on open-vocabulary semantic segmentation.
4.5 Extending to different backbones
To demonstrate the generality of our method, we extend it to several backbones, including ViT-B/16, R50x4, and ConvNeXt-B from OpenCLIP [openclip], as well as ViT-B/16 from EVA-CLIP [sun2023eva]. From Tab. 6, we observe that CLIPSelf [wu2023clipself] is more effective with ViT-based architectures but is less effective with CNN-based architectures, particularly showing degraded performance with R50x4 compared to the baseline. In contrast, our approach yields consistent performance gains across all architectures, thereby broadening its applicability and potential. Furthermore, DenseVLM narrows the performance gap between powerful models like ViT-L/14 fine-tuned by CLIPSelf and lightweight models, with some lightweight models even outperforming the more powerful ones on the Masks-S metric.
4.6 Application to open-vocabulary dense tasks
Experiment settings. To evaluate the performance of the proposed DenseVLM in downstream tasks, we use DenseVLM as the backbone for open-vocabulary dense prediction tasks, including object detection and segmentation. To ensure fairness, our DenseVLM models are trained on the COCO train2017 split with an input resolution of $\times512$.
(a) OV-COCO benchmark
Method Backbone $_{50}^{\mathrm{novel}} AP$$_{50}^{\mathrm{base}}$ AP50
VLDet [lin2022learning] RN50 32.0 50.6 45.8
F-VLM [kuo2022f] RN50 28.0 - 39.6
BARON-Cap [Wu_2023_CVPR] RN50 33.1 54.8 49.1
CORA [CORA] RN50 35.1 35.5 35.4
RO-ViT [Kim_2023_CVPR] ViT-B/16 30.2 - 41.5
RO-ViT [Kim_2023_CVPR] ViT-L/16 33.0 - 47.7
F-ViT+CLIPSelf [wu2023clipself] ViT-B/16 25.4 40.9 36.8
F-ViT+FineCLIP [jing2025fineclip] ViT-B/16 29.8 45.9 41.7
F-ViT [wu2023clipself] ViT-B/16 17.5 41.0 34.9
F-ViT+DenseVLM ViT-B/16 33.1+15.6 52.5+11.5 47.4+12.5
(b) OV-LVIS benchmark
Method Backbone mAPr mAPc mAP
VLDet [lin2022learning] RN50 21.7 29.8 30.1
BARON-Cap [Wu_2023_CVPR] RN50 22.6 27.6 27.6
F-VLM [kuo2022f] RN50 18.6 - 24.2
COR [CORA] RN50x4 22.2 - -
RO-ViT [Kim_2023_CVPR] ViT-B/16 28.0 - 30.2
F-ViT+CLIPSelf [wu2023clipself] ViT-B/16 10.6 7.6 9.3
F-ViT+FineCLIP [jing2025fineclip] ViT-B/16 10.4 8.0 9.5
F-ViT [wu2023clipself] ViT-B/16 11.5 12.3 15.4
F-ViT+DenseVLM ViT-B/16    23.9+12.4 18.4+6.1 21.4+6.0
Table 8: Results on open-vocabulary object detection.
Open-vocabulary semantic segmentation. We apply DenseVLM models initialized with OpenAI CLIP [CLIP] to open-vocabulary semantic segmentation, including SAN [san] with a frozen backbone and Cat-Seg [catseg] with a fine-tuned backbone. The models are trained on COCO-Stuff [caesar2018coco] and evaluated on the ADE20K [ade20k] (ADE-150 and ADE-847) and PASCAL Context [pascal_context] (PC-459) datasets, using the mean Intersection-over-Union (mIoU) metric. As shown in Tab. 7, DenseVLM consistently improves performance across all evaluation datasets, further enhancing the state-of-the-art performance.
Open-vocabulary object detection. Building on previous work [wu2023clipself], we construct open-vocabulary object detectors using the F-ViT architecture, a two-stage detector built on frozen ViTs from EVA-CLIP [sun2023eva]. As shown in Tab. 8, we evaluate performance on the OV-COCO [chen2015microsoft] benchmark by reporting box AP at IoU 0.5 for base, novel, and overall categories ($_{50}^{\mathrm{novel}}, AP$$_{50}^{\mathrm{base}}$, AP50), and on OV-LVIS [Gupta_2019_CVPR] benchmark by mean AP for rare (mAPr), common (mAPc), and all categories (mAP). Replacing the frozen EVA-CLIP ViTs with DenseVLM models leads to clear performance improvements on both benchmarks, achieving competitive performance relative to previous methods.
5 Conclusions
In this paper, we present DenseVLM, a framework designed to mitigate foreground bias in region-level vision-language alignment. DenseVLM can directly replace the original VLMs in open-vocabulary object detection and image segmentation methods, demonstrating consistently clear performance improvements over the baseline models. Furthermore, we validate DenseVLM’s promising scaling ability by exploring efficient region retrieval and decoupled alignment, successfully implementing DenseVLM to the training data from the SA-1B dataset. Overall, DenseVLM offers a generalizable solution for improving dense representations in vision-language models across various backbones, which is essential for open-vocabulary dense prediction tasks.
Acknowledgments
This work was partially funded by NSFC (No. 62225604, 62276145) and the Science and Technology Support Program of Tianjin, China (No. 23JCZDJC01050). This work was also supported by the Supercomputing Center of Nankai University.
[…]
Supplementary Material
We provide an overview of the supplementary materials to ensure a clear and comprehensive understanding.
  • In Sec. 6, we detail the limitations and broader impact of DenseVLM.
  • In Sec. 7, we present the training details.
  • In Sec. 8, we offer supplementary experiments on input image sizes, region proposals, backbones, category sets and threshold of region denoising for training VLMs.
  • In Sec. 9, we show visualizations, including confusion matrices and croppping feayures predictions, to demonstrate foreground bias. .
  • In Sec. 10, we present the dataset information for training and evaluation.
6 Limitations and broader impact
Limitations: Our aim is to develop a region-language alignment model that effectively integrates local visual and semantic features, thereby improving open-vocabulary dense prediction performance. Compared to previous pre-trained Vision-Language Models (VLMs) [openclip, sun2023eva, zhong2022regionclip, wu2023clipself], our proposed DenseVLM achieves superior results and significantly improves downstream task performance. We believe DenseVLM has even greater potential. 1) Scalability. DenseVLM is designed within an efficient, unsupervised region-language alignment framework, making it adaptable to various datasets. However, computational resource limitations have restricted our ability to scale to larger datasets. 2) Model capacity. We employ the ViT-L/14 model from CLIPSelf [wu2023clipself] as a powerful Pre-trained VLM (P-VLM). Utilizing more robust VLMs can yield better performance, and transferring their rich semantic knowledge to training models is a promising direction. 3) Fine-grained Semantic. We categorize objects into broad thing and stuff classes. Fine-grained semantic segmentation and decoupled alignment would enhance the model’s ability to distinguish between similar categories. We plan to explore these avenues in our future research.
Broader impact: DenseVLM exhibits notable potential for open-vocabulary dense predictions within scenes, which can enhance various applications such as robotics and environmental monitoring. By enabling systems to recognize and interpret a wide range of objects and contexts without prior training on specific categories, DenseVLM facilitates more adaptive and versatile applications. Given its broad applicability and non-specialized nature, our method is designed to support a variety of technical advancements without directly addressing specific societal challenges.
item value
image size 512 $\times$ 512
optimizer AdamW [adamw]
learning rate 0.0001
$\beta_{1}$ 0.9
$\beta_{2}$ 0.98
weight decay 0.1
batch size (per card) 48
warmup steps [goyal2018accurate] 1000
epochs 6
learning rate scheduler cosine decay [loshchilov2017sgdr]
number of GPUs 4
automatic mixed precision1 True
Table 9: Training details of DenseVLM.
11footnotetext: https://developer.nvidia.com/automatic-mixed-precision
7 Training details
[…]
For open-vocabulary segmentation, we train the models such as SAN [san] and CAT-Seg [catseg] on the COCO-Stuff [caesar2018coco] dataset for 80kk iterations. For open-vocabulary detection, models are trained for 3 epochs on the OV-COCO [chen2015microsoft] benchmark and 48 epochs on OV-LVIS [Gupta_2019_CVPR] benchmark.
COCO ADE20K
Input GPU Memory Boxes Masks-T Masks-S Boxes Masks-T Masks-S
Image Size (per card) Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
224 9G 60.1 79.9 49.4 62.4 35.3 64.2 40.0 70.0 36.3 56.4 50.3 77.0
320 11G 66.2 85.4 59.2 73.0 41.0 71.2 45.6 76.0 44.0 67.6 54.3 81.7
512 16G 73.4 90.5 71.0 84.8 45.6 77.8 51.3 82.2 52.1 78.0 57.8 85.5
768 27G 74.4 91.3 75.4 90.1 45.5 79.0 52.7 82.9 55.4 82.6 58.2 86.6
1024 39G 76.6 93.1 78.7 93.6 46.5 79.8 53.2 83.6 56.8 83.2 58.6 86.8
Table 10: Ablation study on input image sizes. We report the Top1 and Top5 mean accuracy on classifying boxes and panoptic masks on COCO panoptic and ADE20K panoptic benchmarks. The GPU memory usage corresponds to a batch size of 12 on A40 GPU.
COCO ADE20K
Boxes Masks-T Masks-S Boxes Masks-T Masks-S
Method Region Proposals Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
CLIPSelf 69.1 88.2 66.7 83.0 41.7 75.2 48.1 77.7 47.5 74.2 53.7 82.8
CLIPSelf 70.2 89.2 68.1 83.5 35.7 71.8 49.8 79.7 51.5 76.0 50.9 80.7
DenseVLM 73.4 90.5 71.0 84.8 45.6 77.8 51.3 82.2 52.1 78.0 57.8 85.5
DenseVLM 74.4 91.3 75.4 90.1 45.9 79.0 52.7 82.9 55.4 82.6 58.2 86.6
Table 11: Ablation study on using region proposals. We report the Top1 and Top5 mean accuracy on classifying boxes and panoptic masks (thing and stuff) on COCO panoptic and ADE20K panoptic benchmarks.
8 Additional experiments
Ablation study on input image sizes. To evaluate the effect of input image size on DenseVLM, we conduct experiments with distinct resolutions: 224, 320, 512, 768 and 1024 pixels, for both training and inference. As shown in Tab. 10, model’s performance on the region classification task improves as image resolution increases from 224 to 1024 pixels. This enhancement can be attributed to the greater detail captured at higher resolutions. However, this improvement comes a significant increase in GPU memory usage. Considering the trade-off between computational resources and model performance, we resize the images to $\times512$ pixels to achieve an optimal balance.
Boxes Masks-T Masks-S
$\theta$ Top1 Top5 Top1 Top5 Top1 Top5
0.0 72.1 89.6 68.2 84.3 43.6 76.1
0.1 72.7 90.2 69.1 84.3 44.6 77.2
0.2 73.1 90.4 69.7 84.6 45.1 77.7
0.3 73.4 90.5 71.0 84.8 45.6 77.8
0.4 73.2 90.2 70.2 84.3 45.2 77.5
0.5 73.1 90.0 70.0 84.3 45.0 77.1
0.6 73.1 89.9 69.6 84.0 44.6 76.3
Table 12: Ablation study on threshold of $\theta$ in region denoising.
Ablation study on using region proposals. Following RegionCLIP [zhong2022regionclip] for fine-tuning VLMs with pseudo-labelled region-text pairs, we compare our approach to CLIPSelf [wu2023clipself] in utilizing these pairs. As shown in Tab. 11, CLIPSelf substitutes random image crops with pseudo region-text pairs, resulting in an enhanced recognition for foreground objects while concurrently observing a reduction in the accuracy of background identification. In contrast, our proposed DenseVLM achieves a notable improvement in the recognition accuracy of foreground objects while also improving the identification of background stuff.
Ablation study on the threshold $\theta$. We perform an ablation experiment to assess the impact of varying threshold $\theta$ values of region denoising. As shown in Tab. 12, the model performs the worst when $\theta=0$. When $\theta$ is set lower, the Top-5 accuracy increases, but results in suboptimal performance. This may be due to low-confidence categories causing alignment confusion for the model. Conversely, setting $\theta$ too high filters out too many local images, decreasing performance. By defaul, we select $\theta=0.3$ for DenseVLM.
Ablation study on various backbones. DenseVLM exhibits adaptability to diverse backbones. As shown in Tab. 13, our models achieve consistent superiority over prior approaches [openclip, wu2023clipself] across all dense prediction tasks. Particularly, the ViT-B/16-based DenseVLM performs comparably to the ViT-L/14-based CLIPSelf [wu2023clipself]. Utilizing ViT-L/14 with a large number of parameters as initialization, DenseVLM achieves clearly enhancements across all evaluated metrics, thereby facilitating superior performance in dense prediction tasks.
COCO ADE20K
Boxes Masks-T Masks-S Boxes Masks-T Masks-S
backbones VLMs Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
ViT-B/16 OpenCLIP 49.8 74.3 51.9 72.2 29.2 54.9 28.4 54.1 29.6 53.4 37.9 66.6
ViT-B/16 CLIPSelf 67.6 87.8 64.4 81.2 44.5 77.1 43.4 76.0 44.0 71.1 50.7 82.1
ViT-B/16 DenseVLM* 71.9 90.2 70.0 84.3 47.8 79.4 48.5 79.2 49.0 75.2 55.1 85.2
ViT-B/16 DenseVLM 73.4 90.5 71.0 84.8 45.6 77.8 51.3 82.2 52.1 78.0 57.8 85.5
ViT-L/14 OpenCLIP 21.2 45.3 26.6 48.9 11.2 27.2 48.1 11.9 34.1 13.9 11.1 32.4
ViT-L/14 CLIPSelf 68.3 90.1 67.1 84.5 37.7 71.3 47.1 77.5 47.7 74.4 48.9 82.3
ViT-L/14 DenseVLM* 76.2 92.9 73.3 87.3 47.4 79.1 54.0 84.1 54.2 79.9 57.8 85.9
ViT-L/14 DenseVLM 75.2 91.8 73.3 87.1 45.5 78.1 54.5 85.0 55.6 82.1 58.1 86.4
Table 13: Ablation study on various backbones. We report the Top1 and Top5 mean accuracy on classifying boxes and panoptic masks (thing and stuff) on COCO panoptic and ADE20K panoptic benchmarks. * indicates the model initialized by OpenCLIP [openclip].
COCO ADE20K
Boxes Masks-T Masks-S Boxes Masks-T Masks-S
Categories Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
133 (80) 71.1 88.5 68.7 83.0 44.7 75.2 49.4 79.0 48.5 74.2 54.7 82.8
171 (80) 72.3 89.8 69.4 85.8 44.2 76.0 49.8 79.7 48.9 75.1 55.1 82.4
273 (160) 72.3 89.9 70.1 84.4 44.9 76.4 51.0 81.8 49.3 76.5 57.0 84.0
316 (204) 73.4 90.5 71.0 84.8 45.6 77.8 51.3 82.2 52.1 78.0 57.8 85.5
Table 14: Ablation study on different category sets. We report the Top1 and Top5 mean accuracy on classifying boxes and panoptic masks (thing and stuff) on COCO panoptic and ADE20K panoptic benchmarks.
Ablation study on different category sets. To assess the impact of varying category sets, we conduct experiments using four various sets: 133 (80), 171 (80), 273 (160), and 316 (160), categorized into foreground and background classes. The set of 133 categories exclusively comprises COCO Panoptic [lin2014microsoft] class set, while the 171-category set consists solely of COCO-Stuff [caesar2018coco] class set. The 273-category set integrates non-overlapping classes from both COCO-Stuff and the ADE20K Panoptic [zhou2017scene] dataset, which contains 150 categories. The 316-category set encompasses selected background classes from COCO-Stuff and the ADE20K dataset, which includes 847 categories. Our code accurately reflects all specified category sets.
As shown in Tab. 14, with the increase in the number of categories, the performance of our model progressively improves when evaluated on the COCO Panoptic and ADE20K Panoptic benchmarks. This is because the larger category sets provide a richer representation of objects and stuff, enabling the model to capture more fine-grained information, thereby enhancing its overall performance.
9 Visualizations
Confusion matrix. We compare the region classification results of our method against previous approaches through a confusion matrix visualization for panoptic masks (both thing and stuff categories) on the COCO Panoptic dataset. These confusion matrixes offer a systematic overview of region classification performance, illustrating the incidence of accurate and erroneous classifications, particularly facilitating a precise assessment of models’ accuracy in differentiating between thing and stuff categories. As shown in Fig. 7, Fig. 8, and Fig. 9, prior methods, including EVA-CLIP [sun2023eva], RegionCLIP [zhong2022regionclip], and CLIPSelf [wu2023clipself], often misclassify background regions as co-occurring foreground classes, such as incorrectly identifying snow as skis or sky as kite. In contrast, as demonstrated in Fig. 10, our DenseVLM achieves higher accuracy in recognizing each category, with a notable improvement in the precision of background object identification.
Image grid patches classification. We visualize the classification results of image grid patches using the powerful ViT-L/14 model from CLIPSelf [wu2023clipself]. As shown in Fig. 11, the model focuses heavily on foreground object recognition, but significant portions of background patches are misclassified as foreground objects. The training VLMs are prone to learning these errors. Furthermore, regions with incorrect classifications often have low confidence scores, highlighting the importance of filtering them out.
10 Datasets of training and evaluation
COCO: COCO [lin2014microsoft] is a large-scale panoptic segmentation dataset encompassing 80 Thing and 53 Stuff categories. The dataset comprises 118,000 images designated for the training set and 5,000 images for the validation set.
ADE20k: ADE20k [ade20k] spans a broad spectrum of indoor and outdoor scenes, comprising 2,000 images for the validation set. This dataset includes 100 Thing and 50 Stuff categories. We evaluate open-vocabulary semantic annotations using both the extensive 847-category version (referred to as A-847) and the more frequently adopted 150-category version (referred to as A-150).
Pascal Context: Pascal-Context [everingham2010pascal] constitutes an extensive dataset derived from Pascal-VOC 2010. We evaluate open-vocabulary semantic annotations using the complete set of 459 classes, referred to as PC-459.
OV-COCO: The open-vocabulary detection COCO (OV-COCO) benchmark, introduced in OV-RCNN [Zareian_2021_CVPR], divides the 65 object categories in the COCO dataset into 48 base categories and 17 novel categories.
OV-LVIS: The open-vocabulary detection LVIS (OV-LVIS) benchmark, introduced in ViLD [gu2021open], defines the 337 rare categories from LVIS v1.0 dataset [Gupta_2019_CVPR] as novel categories.
[…]
Figure 7: Confusion matrix visualization for region classification results of EVA-CLIP.
[…]
Figure 8: Confusion matrix visualization for region classification results of RegionCLIP.
[…]
Figure 9: Confusion matrix visualization for region classification results of CLIPSelf.
[…]
Figure 10: Confusion matrix visualization for region classification results of DenseVLM.
[…]
Figure 11: Visualization of image grid patches classification. The powerful ViT-L/14 model exhibits a pronounced focus on foreground object recognition, even when significant portions of background patches are misclassified as foreground objects.